You have two containers of the same liquid. The first container has 98.0 g at T1
Answer #1


apply the principle of calorimetry as

heat lost container 1 = heat lost by container 2

so m1c1 DT =   m2 c2 DT

change of temperature of Contaner 1 is (T-42.6)

change of temperature of container 2 = ( 42.6-21) = 21.6 deg C

so as here sp heat c1   = c2

0.098 * (T1- 42.6) = 0.025 *(21.6)

T1-42.6 = 55.102

T1 = 55.102 + 42.6

T1 = 97.7 deg C --<<<<<<<<<<<<<<<<Answer